please help
solve the following problem.round answer to two decimal places when necessary.
C=100
(Pi)=3.14
D=C÷(Pi)
D=

Answer by stanbon(75887)

solve the following problem.round answer to two decimal places when necessary.
C=100
(Pi)=3.14
D=C÷(Pi)
D= 100/3.14
---
You should use on online calculator for problems of this kind.
Use Google to find "calculator".
---
Ans: 31.85
Cheers,
Stan H.